4x + 8y = 7 3x + 4y = 6
Answer by mananth(16946)

4 x + 8 y = 7 .............1
3 x + 4 y = 6 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -2
4 x 8 y = 7
-6 x + -8 y = -12
Add the two equations
-2 x = -5
/ -2
x = 2.50 OR 5/2
plug value of x in (1)
4 x + 8 y = 7
10 + 8 y = 7
8 y = 7 -10
8 y = -3
y = -0.38 OR -3/8